Find the slope of each line.
The line whose equation is 2x-3y=18
The line whose equation is (1/2)y=-3x+4

Answer by Maths68(1474)

(1)

2x-3y=18
-3y=-2x+18
-3y/-3=(-2x+18)/-3
y=-2x/-3-18/3
y=(2/3)x-6
Compare above equation with the equation of the line
y=mx+b
y-intercept=b=-6
slope of the line=m=2/3


(2)

(1/2)y=-3x+4
y/2=-3x+4
2(y/2)=2(-3x+4)
y=-6x+8
Compare above equation with the equation of the line
y=mx+b
y-intercept=b=8
slope of the line=m=-6